OpenStudy (anonymous):

What is the exact value of the expression √16y^6+9√y^8 − 3√y^6-√y^8 Simplify if possible.

OpenStudy (asnaseer):

ok, the things to remember with these types of problems are:\[\sqrt{ab}=\sqrt{a}\times\sqrt{b}\]\[\sqrt{a^{2n}}=a^n\qquad\text{i.e. square root of an even power just halves the power}\]

OpenStudy (asnaseer):

so, if you look at the first term in your expression, we have:\[\sqrt{16y^6}=\sqrt{16}\times\sqrt{y^6}=4\times y^3=4y^3\]does that make sense?

OpenStudy (anonymous):

@asnaseer yes it does. sorry i was away from my computer

OpenStudy (asnaseer):

now - can you apply the same rules to each of the remaining terms in your expression?

OpenStudy (anonymous):

well you did the first one so 4y^3 9y^4 3y^3 y^4???

OpenStudy (asnaseer):

that is correct, you should now be able to simplify the whole expression.

OpenStudy (anonymous):

once again. thank you so much

OpenStudy (asnaseer):

just gather "like terms"

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

so my final answer should be y^3+8^4 correct?

OpenStudy (asnaseer):

yes - although I think you mis-typed the last term there (missing y)

OpenStudy (anonymous):

yes haha my bad

OpenStudy (asnaseer):

:)